I am 4 years older than someone 14 years younger than someone 6 years older that a 55 year old. How old am I?
Answer by lwsshak3(11628)

I am 4 years older than someone 14 years younger than someone 6 years older than a 55 year old. How old am I?
..
someone 6 years older than a 55 year old=55+6=61
someone 14 younger than the previous someone=61-14=47
I am 4 years older then the previous someone=47+4=51
ans: 51